suppose repetitions are not permitted.
a. how many three digit numbers can be formed from the six digits 2,3,5,7 and 9?
b. how many of these numbers are less than 400?
c. how many are even?
d. how many are odd?
e. how many are multiples of 5?
using combinatorial analysis

Answer by stanbon(75887)

suppose repetitions are not permitted.
a. how many 3-digit numbers can be formed from the six digits 2,3,5,7 and 9?
Ans: 5*4*3 = 60
-----------------------
b. how many of these numbers are less than 400?
1st digit must be 2 or 3.
Ans: 2*4*3 = 24
------------------------
c. how many are even?
Must end in 2.
Ans: 4*3*1 = 12
-------------------------
d. how many are odd?
Cannot end in 2
Ans: 60-12 = 48
----------------------
e. how many are multiples of 5?
Mult end in 5.
Ans: 4*3*1 = 12
=============
Cheers,
Stan H.
------------